OpenStudy (anonymous):

what is the intercept of the line given by the equation below? Enter your answer as an ordered pair

OpenStudy (anonymous):

@CalebBeavers

OpenStudy (anonymous):

@timo86m Is it 8?

OpenStudy (anonymous):

@jim_thompson5910 :)

OpenStudy (anonymous):

Hint y=mx+b

jimthompson5910 (jim_thompson5910):

y = -x+72 is in the form y = mx+b where m = -1 and b = 72 So what do you think the y-intercept is now?

OpenStudy (anonymous):

72? :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

-72

jimthompson5910 (jim_thompson5910):

in general y = mx+b is your line and it has a y-intercept of b ie the point (0, b)

OpenStudy (anonymous):

but dont forget it is in ordered pair

jimthompson5910 (jim_thompson5910):

yes good catch, don't forget that it's really the point (0, 72)

jimthompson5910 (jim_thompson5910):

but you can use shorthand notation to just say 72, depends on the book really

OpenStudy (anonymous):

Oh okay! Lol yeah so All you do really is the intercept B by -1 right? :) @jim_thompson5910 @timo86m

jimthompson5910 (jim_thompson5910):

the slope is -1 because m = -1 compare y = -1x + 72 to y = mx+b and you'll see m and -1 matching, which means m = -1 Similarly, b = 72 because they line up as well